OpenStudy (anonymous):

What is this function written in intercept form? y=-x^2-x+6 A. y=-(x+2)(x+6) B. y=-(x+2)(x-6) C. y=(-x+3)(x-4) D. y=-(x-2)(x+3)

OpenStudy (anonymous):

\(y = -(x^2 + x - 6)\) you want (x^2 + x - 6) to be in the form (x + a)(x + b), expanding the last one: (x + a)(x + b) = x^2 + (a + b)x + ab So what you want to do is find 2 numbers a and b so that a + b = 1, and a * b = -6 then y= -(x +a)(x + b)

OpenStudy (anonymous):

oh okay i see.... so woud it be -2 and 3?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

so d?!

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you so much!